### v1

# MeRF: Motivation-enhanced Reinforcement Finetuning for Large Reasoning Models

Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) has emerged as a powerful learn-to-reason paradigm for Large Language Models (LLMs) to tackle complex reasoning tasks. However, existing RLVR methods overlook one of the most distinctive capabilities of LLMs, their in-context learning ability, as prominently demonstrated by the success of Chain-of-Thought (CoT) prompting. This motivates us to explore how reinforcement learning can be effectively combined with in-context learning to better improve the reasoning capabilities of LLMs. In this paper, we introduce Motivation-enhanced Reinforcement Finetuning (MeRF), an intuitive yet effective method enhancing reinforcement learning of LLMs by involving “telling LLMs the rules of the game”. Specifically, MeRF directly injects the reward specification into the prompt, which serves as an in-context motivation for model to improve its responses with awareness of the optimization objective. This simple modification leverages the in-context learning ability of LLMs aligning generation with optimization, thereby incentivizing the model to generate desired outputs from both inner motivation and external reward. Empirical evaluations on the Knights and Knaves (K&K) logic puzzle benchmark demonstrate that MeRF achieves substantial performance gains over baselines. Moreover, ablation studies show that performance improves with greater consistency between the in-context motivation and the external reward function, while the model also demonstrates an ability to adapt to misleading motivations through reinforcement learning.
Figure 1: Validation Accuracy and Response Length of the RLVR and MeRF models on K&K puzzles benchmark. MeRF shows a significant improvement over RLVR in performance with in-context motivation in the reinforcement finetuning process. Meanwhile, the response length of MeRF models is comparable with the RLVR models, indicating our method of more efficient reasoning capability.
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language understanding and generation tasks, such as instruction following (Ouyang et al., 2022; Zhou et al., 2023; Taori et al., 2023), code generation (Chen et al., 2021; Nijkamp et al., 2022; Zhuo et al., 2024), and medical diagnosis (Singhal et al., 2025; Zhang et al., 2023; Wang et al., 2023). To further improve the reasoning capabilities of LLMs, Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert et al., 2024; Team et al., 2025) has emerged as a promising alternative to conventional supervised fine-tuning approaches (Radford et al., 2018; Brown et al., 2020), as demonstrated by DeepSeek-R1 (Guo et al., 2025) and OpenAI-o1 (Jaech et al., 2024). RLVR treats reasoning as a sequential decision-making process and optimizes models using objective reward signals that can be automatically verified, such as matching mathematical answers to ground truth (Lambert et al., 2024; He et al., 2024), or checking code correctness through unit tests (Le et al., 2022; Liu et al., 2023). This makes RLVR more scalable and robust compared to approaches like Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022; Rafailov et al., 2023), which rely on human preference annotations that can be noisy, inconsistent, and costly to obtain. By optimizing toward clearly defined success criteria, RLVR enables LLMs to iteratively refine their reasoning capabilities, achieving significant performance improvements on complex tasks.
Extending the success of the conventional RL paradigm in gaming, Go, and robotics to LLMs, RLVR relys on external reward signals to guide the training process of LLMs (Zeng et al., 2025; Yu et al., 2025; Hu et al., 2025; Chen et al., 2025). However, RLVR in its current form does not fully exploit the in-context learning capabilities (Wei et al., 2022; Kojima et al., 2022) of LLMs, which is one of their most powerful features. In RLVR, the model receives feedback on the correctness of its outputs, but lacks explicit awareness of the optimization objectives during training. As the verifiable reward function has demonstrated the desired behavior of the model, which can described in natural language, and the in-context learning ability enables LLMs to learn from the given context, an intuitive question is why not tell the LLMs what is the expected behavior, or how is the their output get evaluated, in the reinforcement finetuning process? This is similar to how humans learn: when we are given a task, we often benefit from understanding the rules and objectives before we start working on it. This understanding helps us to align our efforts with the desired outcomes, making the learning process more efficient and effective.
In this paper, we introduce Motivation-enhanced Reinforcement Finetuning (MeRF), a simple yet powerful method that injects the reward specification directly into the prompt, serving as an in-context motivation for model to improve its responses with awareness of the optimization objective. Unlike current RLVR approaches leaving the model blind to the optimization objective during generation, relying on the transcendent reward function to guide the training process, MeRF explicitly informs the model about the reward structure and what constitutes a good response with in-context motivation, incentivizing the model to generate desired outputs from both inner motivation and external reward, leading to more efficient reinforcement finetuning.
Our core contributions are summarized as follows:
  • We propose MeRF, a novel motivation-enhanced reinforcement finetuning method for LLMs, leveraging the in-context learning capabilities of LLMs in the reinforcement finetuning process for efficient improvement of reasoning capabilities.
  • Extensive experiments on the K&K logic puzzle reasoning benchmark show that MeRF significantly outperforms RLVR baseline, validating its effectiveness and efficiency in improving reasoning capabilities of LLMs in complex tasks.
  • We provide comprehensive analysis on the effectiveness of MeRF, including the impact of the consistency between the in-context motivation and the actual reward function, offering insights into the aligned in-context learning with reinforcement finetuning for self-evolving LLMs.
2 Preliminary
2.1 Reinforcement Learning with Verifiable Rewards (RLVR)
Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert et al., 2024) is a reinforcement learning paradigm for traning langauge models on tasks with verifiable outcomes, such as math problems or logic puzzles. The key idea is to use a reward function that can be automatically verified, allowing the model to learn from the ultimate correctness of its outputs. Recent works (Zeng et al., 2025; Yu et al., 2025) have shown that RLVR can significantly improve the reasoning capabilities of LLMs, with reasoning patterns emerged from the optimization for verifiable rewards.
2.2 Group Relative Policy Optimization (GRPO)
Group Relative Policy Optimization (GRPO) (Shao et al., 2024) is a reinforcement learning algorithm designed to optimize policies by leveraging group-wise relative preference information. As a variant of Proximal Policy Optimization (PPO) (Schulman et al., 2017), GRPO foregoes the need for critic models and instead focuses on learning from relative comparisons of actions within groups, significantly enhancing training efficiency for reinforcement finetuning of LLMs. For each question $x$, GRPO samples a group of $G$ outputs $\{y_{i}\}_{i=1}^{G}$ from the policy $\pi_{\theta_{\text{old}}}(\cdot|x)$, and computes the advantage $A_{i}$ for each output $y_{i}$ based on the outcome reward $r_{i}$. The GRPO objective is defined as follows:
$\mathcal{L}_{\text{GRPO}}(\theta)$ $=\mathbb{E}\left[x\sim\mathcal{D};\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x)\right]$ (1)
$=\frac{1}{G}\sum_{i=1}^{G}\left(\min\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\theta_{\text{old}}}(y_{i}|x)}A_{i},\text{clip}\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\theta_{\text{old}}}(y_{i}|x)},1-\varepsilon,1+\varepsilon\right)A_{i}\right)-\beta\mathbb{D}_{KL}\left(\pi_{\theta}||\pi_{\text{ref}}\right)\right),$ (2)
where the advantage
$A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\cdots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\cdots,r_{G}\})}$
is computed by normalizing the outcome rewards of the group of ouputs and sets the advantage for all the tokens in the output $\{y_{i}\}_{i=1}^{G}$. This formulation enables GRPO to learn from relative preferences within each group without the need for a critic model, making it efficient and suitable for the implementation of RLVR to LLMs.
2.3 Knights and Knaves Logic Puzzles
The Knights and Knaves (K&K) (Xie et al., 2024) logic puzzle dataset is a widely used (Xie et al., 2025) benchmark for reinforcement finetuning for LLMs reasoning, which provides a well-structured difficulty level of the logic tasks and allows accurate and easy reward verification for RLVR. The controllable difficulty levels are achieved by varying the number of people in the logic task, the more people in the logic puzzle requiring LLMs more complex reasoning, and more steps to solve the task. The K&K dataset contains 7 different difficulty levels of logic puzzles, with 2 people as the easiest level and 8 people as the most difficult level. Here is an example of the K&K dataset with 3 people:
An Example of K&K Puzzle with 3 People Problem: A very special island is inhabited only by knights and knaves. Knights always tell the truth, and knaves always lie. You meet 3 inhabitants: Penelope, David, and Zoey. Penelope noted, "David is a knight if and only if David is a knave". David told you that Zoey is a knave if and only if Zoey is a knight. According to Zoey, "If Penelope is a knave then David is a knave". So who is a knight and who is a knave? Solution: (1) Penelope is a knave  (2) David is a knave  (3) Zoey is a knight
In this puzzle, the 3 inhabitants are either knights, who always tell the truth, or knaves, who always lie. The statements made by each inhabitant can be analyzed to determine their identities, leading to a unique and verifiable conclusion that Penelope and David are knaves, while Zoey is a knight. The K&K puzzles are challenging logic tasks systematically generated with logic templates, requiring multiple steps of reasoning and logical deduction to arrive at the correct answer. The complexity of the puzzles is precisely controllable by increasing the number of inhabitants. Moreover, the puzzles are unseen in the training of most models, combined with all the above, making it a suitable benchmark for evaluating the reasoning capabilities of RLVR LLMs.
Figure 2: (1) Left: We validate models with motivation and without motivation for inference. Different from the startpoint model, RLVR model achieves a better performance in validation w/ motivation while the motivation is not involved in the training process. We find a connection between the in-context motivation validation and the model optimized through the reinforcement learning process, which is guided by the reward function described in the motivation. (2) Right: The motivation does not directly enhance in-context inference, but instead yields improvements progressively during the RLVR training.
3 Motivation-enhanced Reinforcement Finetuning
In this section, we introduce the Motivation-enhanced Reinforcement Finetuning (MeRF) for efficient LLM reasoning, which enable the LLMs to be aware of the motivation of the reinforcement optimization and motivated by the specific objective in the verifiable reward space.
3.1 Motivation-enhanced Learning
Considering one’s learning process when facing a new task, one may be asked to do something without any information about the task, like "Please solve this math problem." In this case, one may try to solve the problem with the knowledge and experience one has, but it is very likely that one will not be able to solve it correctly, even though actually equipped with enough knowledge for solving it. The reason is that one does not know what is expected and how to do it correctly. One may be able to solve the problem after many trials and errors, but it is very inefficient and time-consuming. Intuitively, it would be much more efficient if one could be told the motivation of the task, like the goal of the task, how to do it correctly, and how would the output likely to be when it is correct. For example, when one is asked to solve a math problem, comparing to be told: "You should think step by step and give the correct answer in the <answer></answer> tag", it will be much more efficient if told: "You should think step by step and give the correct answer in the <answer></answer> tag, if you follow the steps as the demonstration, you will get a score of 1. You will be rewarded with an additional 1 point if you can give the correct answer in the <answer></answer> tag. Otherwise, you will get a score of 0 and no additional points. " This is a clear motivation for humans to solve math problems correctly, providing what is expected and how the metric is for the task.
Recently, RLVR has demonstrated using reinforcement learning to improve the performance of LLMs on various tasks (Team et al., 2025; Guo et al., 2025), expanding the success of reinforcement learning to LLMs while resembling the human learning process. However, the existing reinforcement learning methods for LLMs does not effectively utilize the in-context learning ability of LLMs, which is the key to the success of LLMs. For successful RL training, one of the most important components is a well-defined reward function, which shapes the expected behavior of the policy and guides the training process. In RLVR, the reward function quantifies the correctness of the model’s outputs, and sometimes with other metrics designed by humans for expected behavior. It is true that the reward function designer knows what is the expected behavior of the model, however, will it be possible for LLMs to perform better if they are aware of the motivation of the task, as humans do?
First, we conduct an experiment, comparing the performance of the startpoint model and RLVR model with and without motivation in the inference time. The results are shown in Figure 2, where we find that the startpoint model does not benefit from the motivation in the inference time, while the RLVR model, trained with the reward function described in the motivation, achieves an improvement when validated with in-context motivation. Moreover, we also observe that the motivation does not directly enhance in-context inference, but instead yields improvements progressively during the RLVR training, indicating a connection between the in-context motivation and reinforcement learning process optimized by the transcendent reward function described in motivation. Motivated by this observation, we propose to inject motivation into the training process, leveraging the in-context ability of LLMs for more efficient reinforcement finetuning.
3.2 Motivation with Verified Reward
Following the previous work (Xie et al., 2025) of RLVR, we utilize a verified reward function for the K&K logic puzzle dataset. The reward function contains two components: the Correctness Score and the Format Score. It is implemented by rule-based verification and capable of being described in natural language, enabling the motivation to be injected into the training process. Here is the System Prompt and Motivation for K&K puzzle:
System Prompt and Motivation for K&K Puzzle <|im_start|>system You are a helpful assistant. The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and<answer> </answer> tags, respectively, i.e., <think> reasoning process here </think><answer> answer here </answer>. Now the user asks you to solve a logical reasoning problem. After thinking, when you finally reach a conclusion, clearly state the identity of each character within <answer> </answer> tags. i.e., <answer> (1) Zoey is a knight (2) … </answer>. You will get evaluated following Evaluation Scoring Rules: - Correctness Score: - If your final answer is correct, score 2 - If your answer is understandable but wrong, score -1.5 - If your answer is not parsable or incomplete, score -2 - Format Score: - If you follow the tag format exactly as above, score 1 - Otherwise, score -1 You will get the final score as their sum. Example: (1) The format follows the required structure: +1 (2) The final answer is correct: +2 (3) Total evaluation score: 3 Think carefully, follow the structure, and consider the evaluation rules.<|im_end|>
<|im_start|>user {input the puzzle} <|im_end|>
<|im_start|>assistant <think>
By injecting the motivation into the training process, we enable the model to be aware of the motivation of the task, which describe the reward function of the RLVR pipeline. The motivation provides a clear specification of what is expected and how to do it correctly, aligning the generation with the transcendent optimization objective. This approach leverages the in-context learning ability of LLMs to improve their reasoning capabilities in a more efficient and effective manner.
4 Experiment
In this section, we present the experimental results of our proposed method MeRF for LLM reasoning. We compare the performance of MeRF with the RLVR baseline to demonstrate the effectiveness of motivation-enhanced learning in improving the reasoning capabilities of LLMs in the reinforcement finetuning process.
4.1 Experimental Setup
Models and RL algorithm
We conducted the motivation-enhanced experiments with Qwen2.5 series (Yang et al., 2024) as the startpoint. We used the Qwen2.5-7B-Instruct model as the startpoint model and use GRPO as the RL algorithm for reinforcement finetuning. Qwen2.5-7B-Instruct is a widely used LLM in the research community with a moderate size of 7 billion parameters. Qwen2.5-7B-Instruct is capable of powerful instruct-following for reasoning tasks, which is considered important for reinforcement finetuning of LLMs (Guo et al., 2025). GRPO is an effective and efficient reinforcement learning algorithm for LLMs finetuning, which is suitable for our experiments to demonstrate the effectiveness of MeRF and RLVR baseline in the reinforcement finetuning process without demanding excessive computational resources.
Dataset
We evaluate our method on the K&K logic puzzle dataset. The K&K dataset contains 7 different difficulty levels of logic puzzles, with ranging from 2 people to 8 people in the task. We utilize the 3 to 7 people puzzles for training, the corresponding test set for in-domain evaluation, and the 2 and 8 people puzzles for out-of-distribution (OOD) evaluation. There are 900 samples for training in each difficulty level and 100 samples for evaluation. The total number of samples in the training set is 4500, and 700 samples in the evaluation set.
Other Implementation Details
We implement our method with VeRL (Sheng et al., 2024) and Logic-RL (Xie et al., 2025) codebase. We use the Qwen2.5-7B-Instruct model as the startpoint and conduct the reinforcement finetuning with batch zise of 16, learning rate of 1e-6, 2 epochs of training. We use the rollout framework of Vllm (Kwon et al., 2023) for sampling the outputs of the model. The group size of GRPO is set to 8, clip range is set to 0.2, and the KL penalty is set to 0.001.
Table 1: Performance comparison across models on tasks with varying difficulty by number of people. MeRF demonstrates a significant improvement over the RLVR baseline in both in-domain and OOD scenarios. It is notable all the results are validated without in-context motivation.
Model Difficulty by Number of People Avg.
3 4 5 6 7 Avg. 2 (OOD) 8 (OOD)
o3-mini-high 0.98 0.97 0.95 0.94 0.89 0.95 0.99 0.83 0.94
o1-2024-12-17 0.51 0.38 0.38 0.35 0.30 0.38 0.83 0.20 0.42
Deepseek-R1 0.73 0.77 0.78 0.75 0.88 0.78 0.91 0.83 0.81
GPT-4o 0.57 0.49 0.32 0.23 0.21 0.36 0.68 0.11 0.37
GPT-4o-mini 0.42 0.34 0.17 0.09 0.10 0.22 0.63 0.01 0.25
NuminaMath-7B-CoT 0.13 0.12 0.05 0.01 0.00 0.06 0.28 0.00 0.08
Deepseek-Math-7B 0.21 0.08 0.06 0.02 0.00 0.07 0.35 0.00 0.10
Qwen2.5-7B-Base 0.34 0.16 0.09 0.00 0.00 0.12 0.41 0.00 0.14
Qwen2.5-7B-Instruct 0.24 0.10 0.06 0.04 0.04 0.10 0.43 0.00 0.13
   +RLVR (Baseline) 0.68 0.67 0.57 0.43 0.22 0.51 0.71 0.28 0.51
   +MeRF (Ours) 0.78 0.73 0.68 0.62 0.42 0.65 0.76 0.39 0.63
4.2 Main Results
To demonstrate the effectiveness of MeRF in the reinforcement finetuning process, we compare the performance of MeRF with the RLVR baseline on the K&K logic puzzle dataset. The results in Figure 1 show that MeRF achieves a significant improvement over the RLVR baseline in the validation accuracy within the training process, revealing the remarkable effectiveness of the motivation-enhanced learning for reinforcement finetuning, with only injecting the in-context motivation. In addition, we also observe that the response length of MeRF is comparable with the RLVR models, indicating our method of more efficient reasoning capability.
The results in Table 1 present the validation accuracy in the evaluation of different difficulty levels, comparing the performance of MeRF with the RLVR baseline, startpoint model, and other well-known models. MeRF achieves a significant improvement on the logic reasoning tasks from the startpoint Qwen2.5-7B-Instruct, with only hundreds of training steps, outperforming the RLVR baseline and even some commercial models in all difficulty levels including OOD scenarios. The results of the other baseline models suggest that K&K logic puzzles are challenging tasks for LLMs, and unseen in the training of most models, which further validates the fitness of K&K logic puzzles for analyzing the reasoning capabilities of RLVR models, and the effectiveness of MeRF in the reinforcement finetuning process.
5 Analysis
In this section, we further analyze the effectiveness of MeRF and the impact of motivation on the performance. We conduct the experiments with Qwen2.5-3B-Instruct and Qwen2.5-7B-Instruct as the startpoint models for MeRF and RLVR baseline. We additionaly conduct the experiments of SFT on both models to analyze the performance of SFT with motivation. The results are shown in Table 2.
Table 2: Performance comparison across models on tasks with varying difficulty by number of people.
Model Difficulty by Number of People Avg.
3 4 5 6 7 Avg. 2 (OOD) 8 (OOD)
Qwen2.5-3B-Instruct 0.02 0.00 0.01 0.00 0.00 0.01 0.17 0.00 0.03
   +SFT (w/o motivation) 0.64 0.50 0.25 0.21 0.15 0.35 0.81 0.10 0.38
   +SFT (w/ motivation) 0.28 0.27 0.22 0.18 0.10 0.21 0.34 0.10 0.21
   +RLVR (Baseline) 0.43 0.37 0.20 0.13 0.08 0.24 0.54 0.09 0.26
   +MeRF (Ours) 0.21 0.21 0.19 0.09 0.08 0.16 0.47 0.06 0.19
Qwen2.5-7B-Instruct 0.24 0.10 0.06 0.04 0.04 0.10 0.43 0.00 0.13
   +SFT (w/o motivation) 0.38 0.22 0.14 0.07 0.10 0.18 0.55 0.06 0.22
   +SFT (w/ motivation) 0.54 0.39 0.27 0.13 0.12 0.29 0.60 0.10 0.31
   +RLVR (Baseline) 0.68 0.67 0.57 0.43 0.22 0.51 0.71 0.28 0.51
   +MeRF (Ours) 0.78 0.73 0.68 0.62 0.42 0.65 0.76 0.39 0.63
Figure 3: (1) Left: MeRF underperforms RLVR baseline for Qwen2.5-3B-Instruct. (2) Right: The validation accuracy gap of MeRF is mainly caused by the gap between in-context motivation training and non-motivation validation.
Q1: How does MeRF perform in different models?
As shown in Table 2, the results show that MeRF, on 3B model, underperforms RLVR baseline, while on 7B model, MeRF achieves a significant improvement over RLVR baseline. The validation accuracy of the training curve in Figure 3 (Left) further demonstrates the performance gap between RLVR baseline and MeRF, validating the failure of in-context motivation training for 3B model. In Figure 3 (Right), we observe that when 3B-MeRF is validated with in-context motivation, it achieves comparable performance with 3B-RLVR, and the performance drop is mainly caused by the gap between in-context motivation training and non-motivation validation, shwocasing the 3B model is poor in generalizing to non-motivation validation when trained with in-context motivation. Moreover, we compare it with 7B model in Figure 4, and there is negligible performance drop for 7B model with non-motivation validation, indicating the generalizing capability of model is essential for MeRF to be effective, suggestting that larger models provide the necessary capacity for the method to succeed with MeRF, which is consistent with the findings on the success of RLVR (Yue et al., 2025).
Figure 4: There is negligible performance drop of 7B-MeRF model in non-motivation validation, while 3B-MeRF model fails to generalize to non-motivation validation.
Q2: Can SFT benefit from motivated-enhanced training?
In Table 2, we observe that for 3B model, SFT (w/o motivation) achieves the best performance compared to SFT (w/ motivation), RLVR and MeRF, suggesting that high-quality supervised training with aligned training and validation settings is a more effective strategy for enhancing the performance of small models, which is consistent with the findings of (Guo et al., 2025). The performance drop of the validation gap demonstrated in Figure 3 (Right) also indicates that the model is poor in generalizing to non-motivation validation when SFT with in-context motivation. For 7B model, SFT (w/ motivation) achieves better performance than SFT (w/o motivation), demonstrating that the motivation is beneficial for the performance of SFT, and the model is capable of generalizing to non-motivation validation. However, MeRF achieves the best performance with RLVR baseline outperforming SFT, demonstrating the superiority of reinforcement finetuning for LLMs reasoning (Chu et al., 2025).
Q3: Does the performance improvement come from the in-context inference?
We conduct experiments to investigate where the performance improvement mainly comes from. As results shown in Figure 2 (Left), we observe that for both RLVR and MeRF model, they achieve a slightly better performance in validation with motivation compared to non-motivation validation. However, comparing to the performance improvement of MeRF over RLVR in both non-motivation and motivation validation, we find that the performance improvement of MeRF is mainly from the motivation-enhanced reinforcement finetuning process, demonstrating the effectiveness of the in-context motivation in the training process. The results in Figure 4 also show that the performance of 7B-MeRF does not benefit much from the in-context motivation validation, indicating that the performance improvement of MeRF is not mainly from the in-context inference.
Q4: How do the different motivations affect the performance?
We compare the performance of MeRF with different motivations, including the Motivation (Groud-Truth), which totally matches with the reward function of the optimization process, Motivation (Suboptimal), which is the suboptimal motivation only describing the correctness score, and Motivation (Adverse), which is the adverse motivation misleading the model to provide the wrong answer. The results are shown in Figure 5 (Left) and examples (segments) of the motivations are shown in Figure 5 (Right).
Figure 5: (1) Left: The performance of MeRF with different motivations. The motivation with ground-truth reward function achieves the best performance. Adverse motivation misleads the model to provide the wrong answer, while the model is capable of adapting to the adverse motivation in the process of reinforcement finetuning. (2) Right: Examples of different motivations. Suboptimal motivation only describes the correctness score, while adverse motivation provides the full description of the reward function with opposite scores.
The results demonstrate that the motivation with ground-truth reward function achieves the best performance, and the suboptimal motivation performs better than the RLVR baseline, with additional Correctness Score description included in the motivation. Adverse motivation provides the full description of the reward function as ground-truth motivation, while all the score is assigned to the opposite, which misleads the model to provide the wrong answer. The performance drop of the adverse motivation is mainly caused by the contradiction between the in-context motivation and the reward function of the optimization process. However, after several rounds of unstable learning dynamics, the model adapted to discount the motivation signal, treating the score as uninformative, while the full description of the reward function is still beneficial for the model to outperform the RLVR baseline and Motivation (Suboptimal). The results suggest that MeRF benefits from a better alignment between the motivation and the underlying reward function, resulting in better performance, while the model is capable of adapting to the adverse motivation in the process of reinforcement finetuning.
6 Conclusion
In this paper, we propose a novel method MeRF, leveraging the in-context abilities of LLMs to enhance reinforcment finetuning for reasoning. By injecting the in-context motivation into the training process, MeRF enables the model to be aware of the objective of the task, which aligns the generation with the transcendent optimization objective. Empirical results show that MeRF significantly outperforms RLVR baseline with superior reasoning capabilities. To further understanding the effectiveness of MeRF, we conduct a series of experiments and analysis, revealing the connection between the in-context motivation and the reinforcement learning process optimized by the reward function, providing insights into the potential of motivation-enhanced learning for LLM reasoning and self-evolving LLMs.
Limitations. However, there are still some limitations in our work, presenting the potential for future research. (1) The motivation in MeRF is static in the training process. It is possible to explore the dynamic motivation during the training process in future work. (2) For models of weak generalization capability, how to efficiently implement RLVR and better leverage the in-context motivation for more efficient reinforcement finetuning is still an open question.
Appendix A Appendix
A.1 Broader Impacts
This paper proposes a novel method MeRF for LLM reasoning, leveraging the in-context abilities of LLMs to enhance reinforcement finetuning for reasoning. The proposed method has the potential to improve the performance of LLMs in various reasoning tasks, which can have positive impacts on applications such as natural language understanding, question answering, and logical reasoning. However, it is important to consider how to ensure reliability and safety of the emerged reasoning capabilities of LLMs.
A.2 Related Work
Reinforcement Learning for LLMs. Reinforcement learning (RL) has become a powerful paradigm for fine-tuning LLMs, first demonstrated by reinforcement learning from human feedback (RLHF) to align models with human-preferred responses [Ouyang et al., 2022, Rafailov et al., 2023]. More recently, inspired by the success of DeepSeek-R1 [Guo et al., 2025], Reinforcement Learning with Verified Reward (RLVR) [Lambert et al., 2024] has been proposed to directly enhance reasoning performance by rewarding verifiable success criteria instead of subjective human judgments, which is more similar to the traditional RL paradigm. RLVR has shown promising results in improving the performance of LLMs on various reasoning tasks, including logical reasoning [Xie et al., 2025], coding [Le et al., 2022, Liu et al., 2023], and math problems [Zeng et al., 2025], with simple, designed reward functions. However, these approaches demonstrate the potential of the conventional RL paradigm for LLMs, which may not fully leverage the in-context learning abilities of LLMs, the key to the success of LLMs in previous works [Wei et al., 2022, Brown et al., 2020]. In this work, we propose a novel method MeRF, leveraging the in-context abilities of LLMs to enhance reinforcement finetuning for reasoning.
In-context Learning. In-context learning (ICL) refers to the ability of large language models to learn a task purely from examples and instructions provided in the prompt, without any gradient-based updates to the parameters [Brown et al., 2020]. This ability has been shown to be effective in various tasks, including few-shot learning, zero-shot learning, and even one-shot learning [Wei et al., 2022]. ICL is a key feature of LLMs, enabling them to generalize from a few examples and adapt to new tasks quickly. The success of ICL has led to the development of various prompting strategies, such as few-shot prompting, chain-of-thought prompting, and self-consistency prompting [Wang et al., 2022]. In this work, we investigate the potential of ICL for reinforcement finetuning of LLMs, and propose a novel method MeRF to leverage the in-context abilities of LLMs to enhance reinforcement finetuning for reasoning, injecting the in-context motivation into the training process.
A.3 More Experimental Results
How does MeRF perform in larger models?
To further analyze the performance of MeRF in larger models, we conduct experiments with Qwen2.5-14B-Instruct as the startpoint model. The results are shown in Figure 6, where we observe that MeRF achieves a significant improvement over RLVR baseline in the validation accuracy within the training process, revealing the remarkable effectiveness of the motivation-enhanced learning for reinforcement finetuning, with only injecting the in-context motivation. The results further validate our findings in the Section 5 that larger models provide the necessary capacity for the method to succeed with MeRF, and demonstrate the effectiveness of MeRF in the reinforcement finetuning process.
Case Study.
We provide a case study to illustrate the reasoning capabilities of the final model of MeRF. As shown below, the model is capable of providing the correct answer with a clear reasoning process, understanding the motivation of the task and adhereing to the rules of the metrics. The model is able to analyze the statements of each character and deduce the truthfulness of their claims, leading to the correct conclusion. The model also provides a clear and concise explanation of the reasoning process, demonstrating its ability to understand and apply logical reasoning in a complex scenario.
A Example of Model Response [Puzzle] A very special island is inhabited only by knights and knaves. Knights always tell the truth, and knaves always lie. You meet 3 inhabitants: Evelyn, Benjamin, and William. Evelyn said that If Benjamin is a knight then William is a knight. In a statement by Benjamin: "Evelyn is not a knave". William commented, "Evelyn is a knight". So who is a knight and who is a knave? [Model Response] <think> To solve this, we will analyze each statement based on the rules that knights always tell the truth and knaves always lie. 1. Analyze Evelyn’s statement: "If Benjamin is a knight then William is a knight." - If Evelyn is a knight, Benjamin must be a knight and William must be a knight. (T->T) - If Evelyn is a knave, the statement must be false, so Benjamin being a knight would imply that William is a knave. (F->F) 2. Analyze Benjamin’s statement: "Evelyn is not a knave." - If Benjamin is a knight, his statement must be true, so Evelyn must be a knight. - If Benjamin is a knave, his statement would be false, so Evelyn would be a knave. 3. Analyze William’s statement: "Evelyn is a knight." - If William is a knight, his statement is true, so Evelyn is a knight. - If William is a knave, his statement is false, so Evelyn is a knave. From Benjamin’s statement, if he is a knight, Evelyn must be a knight. From William’s statement, if he is a knight, Evelyn must be a knight. If Evelyn is a knight, for Benjamin’s statement to hold, Benjamin must also be a knight. This configuration satisfies all statements consistently and rules out other scenarios, leading to all three being knights. Therefore, Evelyn, Benjamin, and William are all knights. </think> <answer> (1) Evelyn is a knight (2) Benjamin is a knight (3) William is a knight </answer>
Figure 6: Validation Accuracy and Response Length of Qwen2.5-14B-Instruct with MeRF and RLVR baseline. Similar to the results in Figure 1, MeRF achieves a significant improvement over RLVR baseline in the validation accuracy within the training process, revealing the remarkable effectiveness of the motivation-enhanced learning for reinforcement finetuning, with only injecting the in-context motivation. The results further validate our findings in the Section 5 that larger models provide the necessary capacity for the method to succeed with MeRF.